OpenStudy (anonymous):

Find two positive integers such that their sum is 34 and their product is a maximum.

OpenStudy (anonymous):

17 and 17

OpenStudy (anonymous):

are you supposed to use calculus or pure reason? if calc we make an equation.

OpenStudy (anonymous):

calc

OpenStudy (anonymous):

x+y=34 y=34-x x*(34-x)=maximum = 34x-x² f[x] = 34x-x² maximum happens when df[x]/dx = 0 34-2x = 0 17 = x y = 34-17 = 17

OpenStudy (anonymous):

thank youu